OpenStudy (cherrilyn):

what's the integral of sqrt3 sec^3theta dtheta?

OpenStudy (cherrilyn):

\[\int\limits_{?}^{?} \sqrt{3} \sec ^{3} \theta d \theta\]

OpenStudy (anonymous):

use your log table formula's to express sec differently